Grade 12th passAlgebra

Find the square of the following numbers.
46........

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

(46)2
= (50-4)2
= (50)2 + (4)2 - 2×50×4 [Since, (a+b)2= a2+b2 +2ab]
=2500+16–400
= 2116

Thanks